OpenStudy (lostinmath101):

Find the indicated sum for the following arithmetic series: S18 for 2+9+16+23.....(must show work)

OpenStudy (anonymous):

Do you know the arithmetic series sum formula?

OpenStudy (lostinmath101):

I forgot it.

OpenStudy (anonymous):

You should look it up so you won't forget easily !

OpenStudy (lostinmath101):

Ok... give me a second....

OpenStudy (lostinmath101):

Sn=n(a1+an)/2 \[S _{n}=n(a _{1}+a _{n})\div2\]

OpenStudy (anonymous):

Do you know formula for a_n?

OpenStudy (lostinmath101):

\[a _{n}=a _{1}+(n-1)d\]

OpenStudy (lostinmath101):

So its 2+(18-1)7

OpenStudy (lostinmath101):

Or 121

OpenStudy (anonymous):

yeeeh, you figure it out by yourself :D

OpenStudy (anonymous):

How about Sn?

OpenStudy (lostinmath101):

Is it 1107?

OpenStudy (lostinmath101):

18(2+121)/2

OpenStudy (anonymous):

yep, 9* 123!

OpenStudy (anonymous):

All you do is to give yourself more credit ;)

OpenStudy (lostinmath101):

Thanks so much!!!!

OpenStudy (anonymous):

Remember, without memorize the formula, nobody can help you!

OpenStudy (anonymous):

Of course, it's easily for me to work it out! but I'd be happier if you success in the future w/o my help ;)